OpenStudy (anonymous):

Danny poured 78 liters of water into containers x,y and z. The ratio of the volume of water in contain x to the volume of water in container y is 5:2. The ratio of the volume of water in container y to the volume of water in container z is 8:11. A. how much water was poured into container x. B. how much more water was poured into container x than contain z. This crapy home work is confusing. PLEASE HELPPPPP!!!!!!!!

jigglypuff314 (jigglypuff314):

ratio 5:2 = 20:8 so then x:y:z --> 20:8:11 do you understand that? :)

OpenStudy (anonymous):

yea a little bit

OpenStudy (anonymous):

I get it

jigglypuff314 (jigglypuff314):

then by that ratio 20 + 8 + 11 = 39 but that's not what they want, they want = 78 so 20:8:11 --> 40:16:22 right? then 40 + 16 + 22 = 78 :) ^x ^z then they want to know what is x - z =

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

how much water is poured into container x

jigglypuff314 (jigglypuff314):

40 + 16 + 22 = 78 ^x ^z

jigglypuff314 (jigglypuff314):

so x=40 and z=22 they want to know what is x-z = so what is 40 - 22 = ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but that 22

OpenStudy (anonymous):

b. how much more water was poured into container x than container z .

jigglypuff314 (jigglypuff314):

A. x=40 B. x-z = 40-22=

OpenStudy (anonymous):

no the answer has to be in ratios

jigglypuff314 (jigglypuff314):

then 40:22 to simplify divide both numbers by 2 so 20:11

OpenStudy (anonymous):

ooh ill try

OpenStudy (anonymous):

thankyou